We are doing area and perimeter. The question is, give the dimensions of all the rectangular patios she can make with 64 feet of fencing. Use only integers for dimensions and explain.
Answer by jim_thompson5910(35256)

2L + 2W = P

2L + 2W = 64

L + W = 32 ... divide each term by 2

L = -W + 32

So if W >= 0 and W <= 32, and W is an integer, then you'll get whole number values for L.

Ex: if W = 15, then

L = -W + 32

L = -15 + 32

W = 17

So one possible set of dimensions is 17 by 15